Cats have recently migrated from Malaysia to an uninhibited island in the South China Sea. The gene pool of the migrants includes both alleles for straight (T) and stub (t) tails. You study selection pressures on these cats in the wild, finding the following survivorship in a cohort of 650 kittens.

Phenotype Straight Bent Stub
Genotype TT Tt tt
# kittens 80 414 498
Surviving Adults 70 275 305

1) Calculate survivorship and relative fitness for each genotype

2) Calculate the selection coefficient (s) and degree of dominance (h)
Answer #1

1.Survivorship is survival rate that is the percentage of total born that survive to adult. Survival rate for each genotype -

TT - 70/80×100 =87.5%

Tt - 275/414×100 =66.4%

tt - 305/498×100 =61.2%

Relative fitness(w) for each genotype is survival rate of that genotype divided by maximum survival rate.

w for TT= 87.5/87.5=1

w for Tt= 66.4/87.5=0.76

w for tt= 61.2/87.5=0.699

2. selection coefficient (s) for each genotype is 1- relative fitness of genotype (w)

sTT = 1-1=0

sTt= 1-0.76=0.24

stt=1-0.699=0.30